Two particles P and Q are initially 40m apart P behind Q. P starts moving with uniform velocity of 10m/s towards Q. Q starts from rest and has acceleration 2m/s2 in the dorection of P. Then minimum distance between P and Q will be ?

Grade:11

1 Answers

Vikas TU
7 years ago
From first law of eqn.
v = u + at
10= 0 + 2*t
t = 5s.
In this time
P would have moved 5*10 = 50 m  Q moves 5*5 = 25 m.
THERFEORE, minimum distance  it would cover = >40 + 25 – 50 = 15 m.